Beiträge zur Algebra und Geometrie Contributions to Algebra and Geometry Volume 42 (2001), No. 2, 547-555.

On the Permutation Products of Manifolds

Samet Kera

Institute of Mathematics, P.O.Box 162, 1000 Skopje, Macedonia e-mail: sametasa@iunona.pmf.ukim.edu.mk

Abstract. In this paper it is proven the following conjecture: If G is a subgroup of the permutation group S_{n} and M is a 2-dimensional real manifold, then M^{n}/G is a manifold if and only if G = S_{m_{1}} × S_{m_{2}} × ⋯ × S_{m_{r}} where S_{m_{1}}, …, S_{m_{r}} are permutation groups of partition of {1, 2, …, n} into r subsets with cardinalities m_{1}, …, m_{r}, and M^{n} is the topological product of n copies of M. MSC 2000: 53C99, 20B35 Keywords: permutation products on manifolds, cyclic products on manifolds, permutation group

1. Introduction

Let M be a nonvoid set and let n be a positive integer. In the Cartesian product M^{n} we define a relation ≈ such that (x_{1}, …, x_{n}) ≈ (y_{1}, …, y_{n}) if there exists a permutation ϑ : {1, 2, …, n} → {1, 2, …, n} such that y_{i} = x_{ϑ(i)} (1 ≤ i ≤ n). This is an equivalence relation. The class represented by (x_{1}, …, x_{n}) will be denoted by (x_{1}, …, x_{n})/≈ and the set M^{n}/≈ will be denoted by M^{(n)}. The set M^{(n)} is called a permutation product of M and it was mainly studied in [3].

If M is a topological space, then M^{(n)} is also a topological space. The space (R^{m})^{(n)} (n ≥ 2) is a manifold only for m = 2. If m = 2, then (R^{2})^{(n)} = C^{(n)} is homeomorphic to C^{m}. Indeed, using the fact that the field C is algebraically closed, the mapping φ : C^{(n)} → C^{n} defined by

φ((z_{1}, …, z_{n})/≈) = (σ_{1}, σ_{2}, …, σ_{n})

is a bijection, where σ_{i}(1 ≤ i ≤ n) is the i-th symmetric function of z_{1}, …, z_{n}. The mapping φ is also a homeomorphism. Moreover, if M is a 2-dimensional manifold, then M^{(n)} is a manifold. In [1] it is proven that if M is orientable, i.e. if M is a 1-dimensional complex manifold, then M^{(n)} is a complex manifold. If dimM ≠ 2, then M^{(n)} is not a manifold.

0138-4821/93 $ 2.50 \begin{matrix}c \\ ◯ \end{matrix}2001 Heldermann Verlag

Now let dimM = 2 and let us consider a subgroup G of the permutation group S_{n}. Then define a relation ≈ in M^{n} by

(x_{1}, x_{2}, …, x_{n}) ≈ (x_{τ(1)}, x_{τ(2)}, …, x_{τ(n)})

if and only if τ ∈ G. The factor space M^{n}/≈ will be denoted by M^{n}/G. In [2] it was given the following conjecture which is true for any n ≤ 4.

Conjecture. Let G ≤ S_{n} and M be a 2-dimensional real manifold. Then M^{n}/G is a manifold if and only if G = S_{m_{1}} × S_{m_{2}} × ⋯ × S_{m_{r}} where S_{m_{1}}, …, S_{m_{r}} are permutation groups of a partition of {1, 2, …, n} on r subsets with cardinalities m_{1}, …, m_{r}.

In this paper the conjecture will be proved. We remark that in the special case when G is the cyclic group, then M^{n}/G is called a cyclic product of M and denoted by M^{[n]}. The cyclic product is not a manifold if n > 2 [3] while if n = 2, then it is a manifold which is identical to M^{(2)}. Note also that if dimM ≥ 3, then M^{n}/G is not a manifold whenever G is a cyclic group [3].

2. Proof of the conjecture

Let G be a subgroup of the permutation group S_{n} of the set of n elements {a_{1}, a_{2}, …, a_{n}} which is not of type S_{m_{1}} ×S_{m_{2}} ×⋯×S_{m_{k}} and let M^{n}/G be the corresponding factor-space. We shall prove that M^{n}/G is not a manifold.

The group G defines a partition of the set {a_{1}, a_{2}, …, a_{n}} as follows: if a is an element of {a_{1}, a_{2}, …, a_{n}} then we define

T_{a} = {f(a)| f is a permutation of G}.

Of course, we have that T_{a} and T_{b} are either nonintersecting or equal. Hence they give a partition of the set {a_{1}, a_{2}, …, a_{n}} as claimed.

For these sets T_{a} we define subgroups

G_{a} = {f|f ∈ G and f(x) = x for any x ∉ T_{a}}

and

G′_{a} = {f|f ∈ G and f(T_{a}) = T_{a}}.

It is obvious that G′_{a} = G.

Now we will define some notions and will prove some of their properties.

Definition 1. The cycle of a_{j} with respect to a given permutation f is the finite set {f^{s}(a_{j})|s = 1, 2, …} which is denoted by C(a_{j}, f). The degree of the cycle is the number of its elements.

Each element of the set {a_{1}, a_{2}, …, a_{n}} appears in exactly one cycle with respect to f and any two cycles are nonintersecting or they coincide. If the number of elements of the cycle is 1, then it is called trivial.

Property 1. If p is a prime divisor of the degree of the cycle C(a, f), i.e. if |C(a, f)| = p^{α}m such that (p, m) = 1, then the cycle contains subcycles of degrees p^{i} for i = 1, 2, …, α − 1.

Proof. If s = p^{α−i}m, then the cycle C(a, f^{s}) is a subcycle of degree p^{i} of C(a, f).

Property 2. If C(a, f) and C(b, f) are two nonintersecting cycles with respect to f with degrees p and q, (p, q) = 1, then there exists g ∈ G such that C(a, g) = C(a, f), g = f^{i} for some i ∈ N and C(b, g) = {b} is trivial.

Proof. If we put g = f^{q}, then it is easy to verify that the previous requirements are satisfied.

Property 3. If the degree of C(a_{i}, f) is q > 2 and there exists a permutation h of G_{a_{i}} such that h(c) = d, h(d) = c and h(x) = x for x different from c and d, where c, d ∈ C(a_{i}, f)

– (in other words, h is the transposition (cd) (thereby there exists a positive integer p such that f^{p}(c) = d)) – and if (p, q) = 1, then {g|_{C(a_{i},f)} : g ∈ G_{a_{i}} and g(C(a_{i}, f)) ⊆ C(a_{i}, f)} = {g|_{C(a_{i},f)} : g ∈ G_{a_{i}}} and this is the group of all permutations of the cycle C(a_{i}, f).

Proof. It is sufficient to prove that for each pair of two elements u and v of C(a_{i}, f) there exists h_{1} of G_{a_{i}} such that h_{1}(u) = v and h_{1}(v) = u and for any other element x of C(a_{i}, f), h_{1}(x) = x. It is sufficient to prove that this subgroup contains all transpositions in order to be the group of all permutations. If g = f^{p}, then C(a_{i}, f) = C(a_{i}, g). There exist n_{1}, n_{2} ∈ Z_{q}, n_{1} ≠ n_{2}, such that g^{n_{1}}(u) = d and g^{n_{2}}(v) = d. We can assume that n_{1} < n_{2} since the opposite case can be discussed analogously. Let us define g_{1} = (g∘h)^{n_{2}−n_{1}−1}∘g^{n_{1}}. Then g_{1}(v) = d, g_{1}(u) = c and for any x different from u and v it holds g_{1}(x) = x. Let us define h_{1} = g_{1}^{q−1} ∘ h ∘ g_{1} = g_{1}^{−1} ∘ h ∘ g_{1}, where the second equality holds because g_{1}^{q} = id. Thus we obtain

h_{1}(u) = g_{1}^{−1}(h(g_{1}(u))) = g_{1}^{−1}(h(c)) = v, h_{1}(v) = g_{1}^{−1}(h(g_{1}(v))) = g_{1}^{−1}(h(d)) = u

and

h_{1}(x) = g_{1}^{−1}(h(g_{1}(x))) = g_{1}^{−1}(g_{1}(x)) = x

for any x different from u and v.

Property 4. If the degree q (q > 2) of C(a_{i}, f) is prime number and if there exists a permutation h of G_{a_{i}} such that h(c) = d, h(d) = c and h(x) = x for x different from c and d, where c, d ∈ C(a_{i}, f), then {g|_{C(a_{i},f)} : g ∈ G_{a_{i}}} is the group of all permutations.

Proof. There exists p ∈ N, p < q, such that f^{p}(u) = v. Because q is a prime and p < q, it follows that (p, q) = 1. Now the proof follows from Property 3.

Property 4’. If the degree of C(a_{i}, f) is q and if there exists a permutation h of G_{a_{i}} such that h(c) = d, h(d) = c and h(x) = x for x different from c and d where c, d ∈ C(a_{i}, f) and if f(c) = d where c and d are neighbor with respect to the cycle C(a_{i}, f), then {g|_{C(a_{i},f)} : g ∈ G_{a_{i}}} is the group of all permutations on C(a_{i}, f).

The proof is analogous to that of Property 4. In this case it is p = 1. As a consequence of Properties 3, 4 and 4’, we obtain:

Property 5. If C(a, f) is a cycle with respect to f of degree p > 2 and the transposition (uv) ∈ G for u, v ∈ C(a, f) and (p, s) = p′ > 1, where s is positive integer such that f^{s}(u) = v, then {g|_{C(a_{i},f)} : g ∈ G_{a_{i}}} contains all bijections of the subcycles C(a_{i}, f^{p/p′}), a_{i} ∈ C(a, f), and those obtained by the previous ones by cyclic permutation with f. Furthermore, we have

{g|C(a_{i}, f^{p/p′}) : g ∈ G_{a_{i}}} = {g|C(a_{i}, f^{p/p′}) : g ∈ G_{a_{i}}, g(C(a_{i}, f^{p/p′})) ⊆ C(a_{i}, f^{p/p′})}.

Proof. The row of the cycle C(a_{i}, f^{p/p′}) is equal to p/p′ = q. Since (s, q) = 1 and using Property 3, it follows that

{g|_{C(a_{i},f^{p/p′})} : g ∈ G_{a_{i}}} = {g|_{C(a_{i},f^{p/p′})} : g ∈ G_{a_{i}}, g(C(a_{i}, f^{p/p′})) ⊆ C(a_{i}, f^{p/p′})}

which is the group of all permutations of the cycle (subcycle) C(a_{i}, f^{p/p′}). Each bijection from {g|_{C(a_{i},f^{p/p′})} : g ∈ G_{a_{i}}} belongs also to {g|_{C(a_{i},f)} : g ∈ G_{a_{i}}} and hence it follows that {g|_{C(a_{i},f)} : g ∈ G_{a_{i}}} contains all the bijections of the set of elements of the cycle C(a_{i}, f^{p/p′}), which should be proven.

Note that it may happen to exist f such that C(x, g) = C(x, f) but f^{s} ≠ g for any s.

Definition 2. If C(x, g) = C(x, f) for any x and there exists s such that f^{s}|_{C(x,g)} = g|_{C(x,g)}, then we say that f and g act similarly on their joint cycle C(x, g).

The following question appears naturally. If the degree of the cycle of x with respect to g is minimal as in Property 6, is it possible to find f of G which does not act similarly with g to the cycle C(x, g)? Indeed, the following property holds.

Property 6. If C(x, f) is a cycle which contains x, with the smallest possible degree q, i.e. it does not exist a cycle C(x, g) for g ∈ G of degree smaller than q and bigger than 2, and if C(x, f) = C(x, g), then there exists a positive integer s such that f^{s}|_{C(x,f)} = g|_{C(x,f)}, i.e. f and g act similarly on the cycle C(x, f).

Remark. If q = 2, then it should be g = f. In this case q is a prime number.

Proof. Suppose that g acts on {1, 2, …, q} cyclically. Without loss of generality we can suppose that x = 1 and y = 2, i.e. g(i) = i+1 for i < q and g(q) = 1. Let C(1, f) = C(1, g) and assume that f and g do not act similarly on this common cycle, i.e. there is no number s such that f^{s} = g. Let j be the smallest number from {1, 2, …, q} such that f(j) > g(j), i.e. f(1) = g(1), f(2) = g(2), …, f(j − 1) = g(j − 1), and f(j) > g(j). Therefore it holds g^{q−f(j)}(j) = 1 and q > f(j) − g(j). Let h = g^{q−f(j)} ∘ f^{j−1}. Then it is very easy to verify that h(1) = 1. Indeed, h(1) = g^{q−f(j)} ∘ f^{j−1}(1) = g^{q−f(j)} ∘ g^{j−1}(1) = g^{q−f(j)}(j) = 1. Here we have used that f^{j−1}(1) = j = g^{j−1}(1). From the fact that there is no s such that f^{s} = g, it follows that h is not identity mapping on {1, 2, …, q}. This means that there exists k such that h(k) ≠ k. Hence the cycle C(k, h) has degree smaller than q. This is a contradiction to the choice of q.

Property 6’. If C(a_{i}, f) = C(a_{i}, g), i = 1, 2, …, s, are cycles of prime order q with the property C(a_{i}, f) ∩ C(a_{i}, g) = ∅ for i ≠ j, such that f and g act similarly over each cycle and if f(x) = g(x) = x for any x ∉ C(a_{i}, f), i = 1, …, s, then

M^{n}/<f> = M^{n}/<g> = M^{n}/<f, g> ≅ M^{n−qs} × (M^{s})^{(q)}

where (M^{s})^{(q)} denotes the q-th cyclic product on M^{s}.

Proof. Since f and g act similarly, it follows that f can be generated from g and conversely, i.e. it holds

<f> = <g> = <f, g> ≅ Z_{q}.

If we have in mind how f and g act on M^{n}, we just obtain that

M^{n}/<f> = M^{n}/<g> ≅ M^{n−qs} × (M^{s})^{(q)}.

Property 7. If C(a, f) is a cycle with respect to f with degree m, the transpositions (u_{1}v_{1}) and (u_{2}v_{2}) belong to G such that u_{1}, v_{1}, u_{2}, v_{2} ∈ C(a, f), and s_{1}, s_{2} are divisors of m such that f^{s_{1}}(u_{1}) = v_{1}, f^{s_{2}}(u_{2}) = v_{2}, then the following properties are satisfied:

1. If s_{1} = s_{2}, then <f|_{C(a,f)}, (u_{1}v_{1})> = <f|_{C(a,f)}, (u_{2}v_{2})>.

2. If (s_{1}, s_{2}) = 1, then <f|_{C(a,f)}, (u_{1}v_{1}), (u_{2}v_{2})> ≅ S_{m}.

3. If (s_{1}, s_{2}) = s_{3} > 1, then <f|_{C(a,f)}, (u_{1}v_{1}), (u_{2}v_{2})> ≅ <f|_{C(a,f)}, (u_{3}v_{3})> where u_{3}, v_{3} are elements of the cycle C(a, f) such that f^{s_{3}}(u_{3}) = v_{3} and s_{3} is a divisor of m.

Proof. 1. Using the fact that there exist t, s ∈ N such that f^{t}(u_{1}) = u_{2}, f^{t}(v_{1}) = v_{2} and f^{s}(u_{1}) = v_{1}, we obtain that (u_{2}v_{2}) = f^{t} ∘ (u_{1}v_{1}). From the last equality it follows that <f|_{C(a,f)}, (u_{1}v_{1})> = <f|_{C(a,f)}, (u_{2}v_{2})>.

2. Under the previous assumptions there exist n_{1} and n_{2} ∈ Z_{m} such that n_{2}s_{2} = n_{1}s_{1} + 1 and hence f^{n_{2}s_{2}}(a) = f(f^{n_{1}s_{1}}(a)), i.e. the points f^{n_{2}s_{2}}(a) and f^{n_{1}s_{1}}(a) are neighbors in the cycle C(a, f). Moreover we will prove that the transpositions (af^{n_{2}s_{2}}(a)) and (af^{n_{1}s_{1}}(a)) belong to G. Indeed, there exists s such that f^{s}(a) = u_{1}, f^{s_{1}}(f^{s}(a)) = f^{s_{1}}(u_{1}) = v_{1} = f^{s}(f^{s_{1}}(a)) and hence we obtain that (u_{1}v_{1}) = f^{s} ∘ (af^{s_{1}}(a)) ∈ G. Since G is a group, it follows that (af^{s_{1}}(a)) ∈ G. Note that the point a is an arbitrary point from the cycle C(a, f) and hence (f^{t}(a)f^{t}(f^{s_{1}}(a))) ∈ G. Finally, we obtain

(af^{n_{1}s_{1}}(a)) = (af^{s_{1}}(a)) ∘ (f^{s_{1}}(a)f^{2s_{1}}(a)) ∘ (f^{2s_{1}}(a)f^{3s_{1}}(a)) ∘ ⋯
⋯ ∘ (f^{(n_{1}−1)s_{1}}(a)f^{n_{1}s_{1}}(a)) ∘ (f^{(n_{1}−1)s_{1}}(a)f^{(n_{1}−2)s_{1}}(a)) ∘ ⋯
⋯ ∘ (f^{2s_{1}}(a)f^{s_{1}}(a)) ∘ (f^{s_{1}}(a)a) ∈ G.

Analogously, it verifies that (af^{n_{2}s_{2}}(a)) ∈ G. Thus we have

(f^{n_{2}s_{2}}(a)f^{n_{1}s_{1}}(a)) = (af^{n_{2}s_{2}}(a))(af^{n_{1}s_{1}}(a))(af^{n_{2}s_{2}}(a)) ∈ G.

Now using the Property 4’, we obtain that each bijection of the cycle C(a, f) is generated. This completes the proof of 2.

3. Analogously there exist n_{1} and n_{2} ∈ Z_{m} such that n_{1}s_{1} +s_{3} = n_{2}s_{2} and hence it follows that

f^{n_{2}s_{2}}(a) = f^{s_{3}}(f^{n_{1}s_{1}}(a)).

Suppose that f^{n_{2}s_{2}}(a) = a_{2} and f^{n_{1}s_{1}}(a) = a_{1}. The points a_{1} and a_{2} belong to C(a, f), hence we have

(a_{1}a_{2}) = (aa_{1}) ∘ (a_{1}a_{2}) ∘ (aa_{2}) ∈ G

and f^{s_{3}}(a_{2}) = a_{1}. The transpositions (af^{s_{2}}(a)) and (af^{s_{1}}(a)) can be obtained by composing transpositions of the form (bf^{s_{3}}(b)). Moreover, for the transposition (af^{s_{2}}(a)) it holds s_{2} = ps_{3} for a positive integer p. Hence it holds that

\begin{aligned}(af^{s_{2}}(a)) = & (af^{s_{3}}(a)) ∘ (f^{s_{3}}(a)f^{2s_{3}}(a)) ∘ ⋯ ∘ (f^{(p−1)s_{3}}(a)f^{ps_{3}}(a))∘ \\ & ∘ (f^{(p−1)s_{3}}(a)f^{(p−2)s_{3}}(a)) ∘ (f^{(p−2)s_{3}}(a)f^{(p−3)s_{3}}(a)) ∘ ⋯ ∘ (f^{s_{3}}(a)a) \end{aligned}

and analogously for (af^{s_{1}}(a)). This means that each bijection which can be generated by f|_{C(a,f)}, (af^{s_{2}}(a)) and (af^{s_{1}}(a)) can also be generated only by f|_{C(a,f)} and (af^{s_{3}}(a)). The converse holds, too. From (af^{s_{2}}(a)) ∈ G and from the fact that a and f^{s_{i}}(a) are neighbors with respect to the cycle C(a, f^{s_{i}}) according to Property 4’, it follows that each bijection restricted on the cycle C(a, f^{s_{i}}) belongs to G, and hence also the transposition (af^{n_{i}s_{i}}(a)) belongs to G for i = 1, 2. By substituting f^{n_{1}s_{1}}(a) instead of a we get

(f^{n_{1}s_{1}}(a)f^{n_{2}s_{2}}(f^{n_{1}s_{1}}(a))) ∈ G,
(f^{n_{1}s_{1}}(a)f^{s_{3}}(a)) ∈ G,
(a, f^{s_{3}}(a)) = (af^{n_{1}s_{1}}(a)) ∘ (f^{n_{1}s_{1}}(a)f^{s_{3}}(a)) ∘ (af^{n_{1}s_{1}}(a)) ∈ G

and hence 3. is proven.

Remark. From Property 7 we obtain the following conclusion. If G restricted on the cycle C(a, f) does not contain all its bijections, then all bijections restricted on the cycle C(a, f) obtained by G, can be obtained by f and the transposition (uv) (if such exists) for u, v ∈ C(a, f) such that s|p and (s, p) > 1 is the smallest one with that property, where s is the smallest positive integer which satisfies f^{s}(u) = v.

Now let us return to the proof of the conjecture. Note that G acts transitively on T_{a} for any a ∈ {a_{1}, …, a_{n}}. Moreover, since G is not a group of the form S_{m_{1}}×S_{m_{2}}×⋯×S_{m_{k}}, at least one of the subgroups G_{a} or G′′_{a} is not of the form S_{m_{i}}, where G′′_{a} = G|_{T_{a}} = {f|_{T_{a}} : f ∈ G}. Thus we have two cases:

1. There exists g such that there are at least two nonintersecting minimal cycles C(a, g) and C(b, g) with the same degree q (here ”minimal” is meant in the sense of Property 6 even if the degree can be 2 in this case), on which g acts simultaneously, and thereby there is no h ∈ G such that h(x) = x for any x ∉ C(a, g), and h(x) ≠ x for any x ∈ C(a, g). Indeed, h moves only the points of the cycle C(a, g) and thereby the cycle C(a, h) is a subcycle of C(a, g).

2. There is no g ∈ G with the previous property. In both cases we will prove that M^{n}/G is not a manifold, where M is a 2-dimensional manifold. Case 1. Assume that the degree of the cycles C(a, g) and C(b, g) is 2. There exist u, v, x, y such that the composition (uv)(xy) enters in the decomposition of g but the transpositions (uv) and (xy) are not elements of G. Let g be chosen such that the number of its cycles of degree 2 over which g acts simultaneously is r > 1. The number r will be called pairwise degree of g. Let g be chosen with the smallest possible value of r.

If we choose a point x = (x_{1}, x_{2}, …, x_{n}) ∈ M^{n} such that the coordinates corresponding to the same cycle of degree 2 of g are equal, i.e. x_{g(i)} = x_{i} for all indices i with the property g^{2}(i) = i, then the points of different cycles are different and the remaining points are completely different. Here g is a bijection on the index set {1, 2, …, n}.

The set G^{g} = {f ∈ G| f acts invariantly on the point x} = {f ∈ G|f acts invariantly on any cycle of g} is a group. Therefore, the minimality of r implies that if g and f have the same pairwise degree r, then f = g. Hence we obtain that G^{g} = {id, g}.

Now we note that the tangent space at the point corresponding to x in the factor space is homeomorphic to

(R^{2})^{n}/G^{g} ≅ ((R^{2})^{r})^{[2]} × R^{2} × ⋯ × R^{2}

which is not homeomorphic to R^{2n} and hence the space M^{n}/G is not a manifold.

If the cycles C(a, g) and C(b, g) have degree q > 2 with the previous property and if we have in mind Properties 1 and 2, then C(a, g) and C(b, g) can be chosen such that q is a prime number. Further, from Property 4 we can conclude that there is no subcycle of degree 2, i.e. transposition for none of the previous cycles. Otherwise, Property 4 would imply that any bijection on the corresponding cycle could be obtained and hence g would not satisfy the conditions from 1, i.e. there does exist h as in 1.

Further, let us assume that g is chosen such that there exists the smallest possible number r > 1 for nonintersecting cycles on which g acts simultaneously as above.

We choose a point x = (x_{1}, x_{2}, …, x_{n}) ∈ M^{n} such that all coordinates corresponding to the same cycle of the previous r cycles of g are equal. There the points of different cycles are different and the remaining points are completely different. In this case the tangent space at the point x^{≈} ∈ M^{n}/G which corresponds to x is homeomorphic to (R^{2})^{n}/G^{g}, where G^{g} = {f|f(C(x′, g)) = C(x′, g) for any cycle C(x′, g) of g}. By the minimality of q and r and from Property 6, it follows that G^{g} = <g> = Z_{r}, since any f ∈ G^{g} acts similarly to g on any cycle of g, i.e. over any such cycle f^{s} = g restricted on it, for some positive integer s. Thus we get that

(R^{2})^{n}/G^{g} ≅ ((R^{2})^{r})^{[q]} × R^{2} × ⋯ × R^{2}

where R^{2} appears n − qr times. This space is not homeomorphic to R^{2n} and thus we obtain that M^{n}/G is not a manifold.

Case 2. In this case for any cycle C(x, g) there exists f ∈ G such that C(x, g) = C(x, f) and thereby f(y) = y for any y ∉ C(x, f), i.e. any cycle can be considered separately and the cycle C(x, f) is called unical. Because of this argument we obtain that G_{a} = G for any a. Further we will need the following property.

Property 8. Under the assumptions in case 2, if for any unical cycle obtained from G any bijection on that cycle is contained in G, then G_{a} = G ≅ S_{m_{i}}, where m_{i} = |T_{a}|.

Proof. Since G acts transitively on T_{a}, it follows that for any u, v ∈ T_{a} there exist f and g ∈ G such that g(u) = v and C(u, g) = C(u, f), and thereby the cycle C(u, f) is unical. But since v ∈ C(u, f), there exists a positive integer s such that f^{s}(u) = v. According to the assumption that any bijection on the cycle C(a, f) belongs to G, we obtain that also the transposition (uv) belongs to G. Since u and v are arbitrary elements of the set T_{a}, it follows that any bijection in T_{a} can be generated because f acts trivially on the elements not in T_{a}.

Thus there exists at least one cycle C(a, f) as above for which not all transpositions (uv); u, v ∈ C(a, f), belong to G. We choose G(a, f) with degree p > 2 which is minimal with this property. Note that in this case p may not be prime. According to Properties 3 and 5, there are two possibilities:

a) The cycle C(a, f) does not contain u, v such that the transposition (uv) belongs to G.

b) Suppose the transposition (uv) belongs to G for some u, v ∈ C(a, f). If s is the smallest positive integer such that f^{s}(u) = v, then (p, s) = p_{1} > 1. We consider now both of these possibilities.

a) From the minimality of the degree of the cycle C(a, f) and from Property 6, we obtain that for any g ∈ G such that C(a, f) = C(a, g), there is a positive integer s such that g = f^{s}, i.e. f and g act similarly on C(a, f).

We choose a point x = (x_{1}, …, x_{n}) ∈ M^{n} such that all coordinates corresponding to the previous cycle of f are equal. There the points of different cycles are different and the remaining points are completely different.

In this case the tangent space at the point x^{≈} ∈ M^{n}/G which corresponds to x is homeomorphic to (R^{2})^{n}/G^{f}, where G^{f} = {g|g(C(a, f)) = C(a, f)}. Using Property 6 and the minimality of p, it follows that G^{f} = <f> = Z_{p}, which implies that M^{n}/G^{f} is homeomorphic to (R^{2})^{[p]} × R^{2} × ⋯ × R^{2}. Since this space is not homeomorphic to R^{2n} for p > 2, the factor space M^{n}/G is not a manifold.

b) In this case since G is not of the form S_{m_{1}} × S_{m_{2}} × ⋯ × S_{m_{k}} it follows that there exists a cycle C(a, f) such that f contains only the cycle C(a, f) as non-trivial and thereby G does not contain all permutations of that cycle. According to Property 7, there exists a transposition (uv) ∈ G, u, v ∈ C(a, f) such that for any g ∈ G such that g(C(a, f)) = C(a, f) is generated by f and (uv), and thereby f^{s}(u) = v, for which (s, m) = p_{1} > 1 and p_{1} < p. Thus G|_{C(a,f)} is isomorphic to <f, (uv)>.

Thus we consider a point x = (x_{1}, …, x_{n}) ∈ M^{n} where the coordinates corresponding to the points of the cycle C(a, f) are equal. There the points of different cycles are different and the remaining points are completely different. Then the tangent space over the corresponding point of M^{n}/G is homeomorphic to (R^{2})^{n}/G^{f}, where G^{f} = {g ∈ G|g(C(a, f)) = C(a, f)} ≅ G|_{C(a,f)} ≅ <f, (uv)>. But (R^{2})^{n}/G^{f} is homeomorphic to ((R^{2})^{(p_{1})})^{[p/p_{1}]} × R^{2} × ⋯ × R^{2}. Since p_{1} > 1 and p/p_{1} > 1, this space is not a manifold. Thus M^{n}/G is not a manifold.

References

[1] Trenčevski, K.: Permutation products of 1-dimensional complex manifolds. Contributions, Mac. Acad. Sci. Arts. XX (1999), in press.

[2] Trenčevski, K.; Kera, S.: One conjecture concerning the permutation products on manifolds. Mathematica Balkanica 12 (1998), 425–429.

[3] Wagner, C. H.: Symmetric, cyclic and permutation products of manifolds. Dissert. Math. (Rozprawy Mat.) 182 (1980), 3–48.

Received October 25, 1999; revised version September 12, 2000